OpenStudy (anonymous):

In 1991 the life expectancy of males in a certain country was 66.3 years. In 1996 it was 69.6 years. Let E represent the life expectancy in year t and let t represent the number of years since 1991. The linear function E(t) that fits the data is????? E(t)=?+? E(14)=?

OpenStudy (saifoo.khan):

i will pass this one.

OpenStudy (anonymous):

lol thanks

OpenStudy (saifoo.khan):

haha, i always leave this blank

OpenStudy (anonymous):

lol me too

OpenStudy (anonymous):

E(t)=33/50t+66.3 E(14)=75.54

OpenStudy (saifoo.khan):

how you got 33/50?

OpenStudy (anonymous):

finding the slope between the two given points (1991,66.3) and (1996,69.6) is 3.3/5 or 33/50 the age at t=0 is 66.3

OpenStudy (saifoo.khan):

i was always confused about it's slope thanks man!